# Improved Off-policy Reinforcement Learning in Biological Sequence Design

### v1

@@ Abstract
Designing biological sequences with desired properties is a significant challenge due to the combinatorially vast search space and the high cost of evaluating each candidate sequence. To address these challenges, reinforcement learning (RL) methods, such as GFlowNets, utilize proxy models for rapid reward evaluation and annotated data for policy training. Although these approaches have shown promise in generating diverse and novel sequences, the limited training data relative to the vast search space often leads to the misspecification of proxy for out-of-distribution inputs. We introduce $\delta$-Conservative Search, a novel off-policy search method for training GFlowNets designed to improve robustness against proxy misspecification. The key idea is to incorporate conservativeness, controlled by parameter $\delta$, to constrain the search to reliable regions. Specifically, we inject noise into high-score offline sequences by randomly masking tokens with a Bernoulli distribution of parameter $\delta$ and then denoise masked tokens using the GFlowNet policy. Additionally, $\delta$ is adaptively adjusted based on the uncertainty of the proxy model for each data point. This enables the reflection of proxy uncertainty to determine the level of conservativeness. Experimental results demonstrate that our method consistently outperforms existing machine learning methods in discovering high-score sequences across diverse tasks—including DNA, RNA, protein, and peptide design—especially in large-scale scenarios. The code is available at https://github.com/hyeonahkimm/delta_cs.
@@ 1 Introduction
Designing biological sequences with desired properties is crucial in therapeutics and biotechnology (Zimmer, 2002; Lorenz et al., 2011; Barrera et al., 2016; Sample et al., 2019; Ogden et al., 2019). However, this task is challenging due to the combinatorially large search space and the expensive and black-box nature of objective functions. Recent advances in deep learning methods for biological sequence design have shown significant promise at overcoming these challenges (Brookes & Listgarten, 2018; Brookes et al., 2019; Angermueller et al., 2020; Jain et al., ).
Among various approaches, reinforcement learning (RL), which leverages a proxy model as a reward function, has emerged as one of the successful paradigms for automatic biological sequence design (Angermueller et al., 2020). RL methods have the benefit of exploring diverse sequence spaces by generating sequences token-by-token from scratch, enabling the discovery of novel sequences. They employ deep neural networks as inexpensive proxy models to approximate costly oracle objective functions. The proxy model serves as a reward function for training deep RL algorithms, enabling the policy network to generate high-reward biological sequences. The model can be trained in an active learning manner (Gal et al., 2017), iteratively annotating new data by querying the oracle with points generated using the policy; these iterations are called query rounds. There are two approaches for training the policy in this context: on-policy and off-policy.
@@
DyNA PPO (Angermueller et al., 2020), a representative on-policy RL method for biological sequence design, employs Proximal Policy Optimization (PPO; Schulman et al., 2017) within a proxy-based active training loop. While DyNA PPO has demonstrated effectiveness in various biological sequence design tasks, its major limitation is the limited search flexibility inherent to on-policy methods. It cannot effectively leverage off-policy data points, such as data collected from previous rounds.
@@
Conversely, Generative Flow Networks (GFlowNets; Bengio et al., 2021), off-policy RL methods akin to maximum entropy policies (Tiapkin et al., 2024; Deleu et al., 2024), offer diversity-seeking capabilities and flexible exploration strategies. Jain et al. applied GFlowNets to biological sequence design with additional Bayesian active learning schemes. They leveraged the off-policy nature of GFlowNets by mixing offline datasets with on-policy data during training. This approach provided more stable training compared to DyNA PPO and resulted in better performance.
@@
However, recent studies have consistently reported that GFlowNets perform poorly on long-sequence tasks such as green fluorescent protein (GFP) design (Kim et al., 2023; Surana et al., 2024). We hypothesize that this poor performance stems from the insufficient quality of the proxy model in the early rounds. For example, in the typical benchmark, the proxy is trained with 5,000 sequences, while GFPs have a combinatorial search space of $20^{238}>10^{309}$. GFlowNets are capable of generating sequences from scratch and producing novel sequences beyond the data points. However, when the quality of the proxy model is unreliable, novel sequences that are out-of-distribution to the proxy model yield unreliable results (Trabucco et al., 2021; Yu et al., 2021). This motivates us to introduce a conservative search strategy to restrict the trajectories to the neighborhood of the existing data point.
@@
Contribution. In this paper, we propose a novel off-policy search method called $\delta$-Conservative Search ($\delta$-CS), which enables a trade-off between sequence novelty and robustness to proxy misspecification by using a conservativeness parameter $\delta$. Specifically, we iteratively train a GFlowNet using $\delta$-CS as follows: (1) we inject noise by independently masking tokens in high-score offline sequences with Bernoulli distribution with parameter $\delta$; (2) the GFlowNet policy sequentially denoises the masked tokens; (3) we use these denoised sequences to train the policy. When $\delta=1$ (full on-policy), this becomes equivalent to sampling full trajectories, and when $\delta=0$ (fully conservative), this reduces to only showing offline sequences in off-policy training. We adaptively adjust $\delta(x;\sigma)$ using the proxy model’s uncertainty estimates $\sigma(x)$ for each data point $x$. This approach allows the level of conservativeness to be adaptively adjusted based on the prediction uncertainty of each data point. Figure 1 illustrates the overall procedure of the $\delta$-CS algorithm.
@@
Our extensive experiments demonstrate that $\delta$-CS significantly improves GFlowNets, successfully discovering higher-score sequences compared to existing model-based optimization methods on diverse tasks, including DNA, RNA, protein, and peptide design. This result offers a robust and scalable framework for advancing research and applications in biotechnology and synthetic biology.
@@ 2 Problem Formulation
We aim to discover sequences $x\in\mathcal{V}^{L}$ that exhibit desired properties, where $\mathcal{V}$ denotes the vocabulary, such as amino acids or nucleotides, and $L$ represents the sequence length, which is usually fixed. The desired properties are evaluated by a black-box oracle function $f:\mathcal{V}^{L}\to\mathbb{R}$, which evaluates the desired property of a given sequence, such as binding affinity or enzymatic activity. Evaluating $f$ is often both expensive and time-consuming since it typically involves wet-lab experiments or high-fidelity simulations.
Advancements in experimental techniques have enabled the parallel synthesis and evaluation of sequences in batches. Therefore, lab-in-the-loop processes are emerging as practical settings that enable active learning. Following this paradigm, we perform $T$ rounds of batch optimization, where in each round, we have the opportunity to query $B$ batched sequences to the (assumed) oracle objective function $f$. Due to the labor-intensive nature of these experiments, $T$ is typically very small. Following Angermueller et al. (2020) and Jain et al. , we assume the availability of an initial offline dataset $\mathcal{D}_{0}=\{(x^{(n)},y^{(n)})\}_{n=1}^{N_{0}}$, where $y=f(x)$. The initial number of data points $N_{0}$ is typically many orders of magnitude smaller than the size of the search space, as mentioned in the introduction. The goal is to discover, after $T$ rounds, a set of sequences that are novel, diverse, and have high oracle function values.
@@ 3 Active Learning for Biological Sequence Design
Following Jain et al. , we formulate an active learning process constrained by a budget of $T$ rounds with query size $B$, is executed through an iterative procedure consisting of three stages with a novel component of $\delta$-Conservative Search ($\delta$-CS) which will be detailed described in Section 4:
@@ 3.1 Step A: Proxy Training
Following Jain et al. , we train the proxy model $f_{\phi}$ using the dataset $\mathcal{D}_{t-1}$ by minimizing the mean squared error loss:
@@ 3.2 Step B: Policy training with $\delta$-CS
For policy training, we employ GFlowNets, which aim to produce samples from a generative policy where the probability of generating a sequence $x$ is proportional to its reward, i.e.,
$p(x;\theta)\propto R(x;\phi)=f_{\phi}(x)+\kappa\sigma(x).$ (2)
Following Jain et al. , the reward $R(x;\phi)$ is defined as $f_{\phi}(x)+\kappa\sigma(x)$, which combines the proxy value $f_{\phi}(x)$ and the uncertainty $\sigma(x)$ in the form of the upper confidence bound (UCB; Srinivas et al., 2010) acquisition function. This approach prioritizes regions with higher uncertainty, enabling us to query them in the next active round. Here, $\kappa$ is a mixing hyperparameter.
Policy parameterization. The forward policy $P_{F}$ generates state transitions sequentially through trajectories $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{L}=x)$, where $s_{0}=()$ represents the empty sequence, and each state transition involves adding a sequence token. The full sequence $s_{L}=x$ is obtained after $L$ steps, where $L$ is the length of the sequences. The forward policy $P_{F}(\tau;\theta)$ is a compositional policy defined as:
@@
Learning objective and training trajectories. The policy is trained to minimize TB loss: $\mathcal{L}_{\text{TB}}(\tau;\theta)$:
@@
Usually, GFlowNets training is employed to minimize TB loss with training trajectories $\tau$ on full supports, asymptotically guaranteeing optimality for the distribution:
$p(x;\theta)\propto R(x;\phi).$
A key challenge in prior works Jain et al. is that the proxy model $f_{\phi}(x)$ often produces highly unreliable rewards $R(x;\phi)$ for out-of-distribution inputs. In our approach, we mitigate this by providing off-policy trajectories within more reliable regions by injecting conservativeness into off-policy search. Therefore, we minimize TB loss with $\delta$-CS, which offers controllable conservativeness.
4 $\delta$-CS: controllable conservativeness in off-policy search
This section introduces $\delta$-Conservative Search ($\delta$-CS), an off-policy search method that enables controllable exploration through a conservative parameter $\delta$. The parameter $\delta$ defines the Bernoulli distribution governing the masking of tokens in a sequence. Our algorithm is iterations of the following:
@@ 4 $\delta$-CS: controllable conservativeness in off-policy search
Rank-based reweighted prior. First, we sample a reference sequence $x$ from the prior distribution $P_{\mathcal{D}_{t-1}}$. To exploit high-scoring sequences, we employ rank-based prioritization (Tripp et al., 2020).
@@
$P_{\text{noise}}(\tilde{x}\mid x,\delta)=\prod_{i=1}^{L}\left[\delta\cdot\mathbb{I}\{\tilde{e}_{i}=\text{[MASK]}\}+(1-\delta)\cdot\mathbb{I}\{\tilde{e}_{i}=e_{i}\}\right],$
@@
$P_{\text{denoise}}(\hat{e}_{t}\mid\hat{s}_{t-1},\tilde{x};\theta)=\begin{cases}\mathbb{I}\{\hat{e}_{t}=\tilde{e}_{t}\},&\text{if }\tilde{e}_{t}\neq\text{[MASK]},\\ P_{F}(\hat{s}_{t}=(\hat{s}_{t-1},\hat{e}_{t})\mid\hat{s}_{t-1};\theta),&\text{if }\tilde{e}_{t}=\text{[MASK]}.\end{cases}$
@@
Algorithm 1 Active Learning GFlowNets with $\delta$-CS
2:procedure $\delta$-CS ($\mathcal{D}_{t-1},M,\delta$) $\triangleright$ $\delta$-CS subroutine
3:     sample high reward data $x_{1},\ldots,x_{M}$ with rank-based reweighed prior $P_{\mathcal{D}_{t-1}}(\cdot)$.
4:     obtain masked data $\tilde{x}_{1},\ldots,\tilde{x}_{M}$ with noise injection policy $P_{\text{noise}}(\tilde{\cdot}|\cdot,\delta)$ from $x_{1},\ldots,x_{M}$.
5:     obtain denoised $\hat{x}_{1},\ldots,\hat{x}_{M}$ with denoising policy $P_{\text{denoise}}(\hat{\cdot}\mid\tilde{\cdot};\theta)$ from $\tilde{x}_{1},\ldots,\tilde{x}_{M}$.
6:     return $\hat{x}_{1},\ldots,\hat{x}_{M}$.
7:end procedure
8:for $t=1$ to $T$ do $\triangleright$ Active learning with $T$ rounds
9:     while proxy training iterations do $\triangleright$ Step A: Proxy training
10:         train proxy $f_{\phi}(x)$ with current round dataset $\mathcal{D}_{t-1}$:
$\mathcal{L}(\phi)=\mathbb{E}_{x\sim P_{\mathcal{D}_{t-1}}(x)}\left[\left(f(x)-f_{\phi}(x)\right)^{2}\right].$
11:     end while
12:     while policy training iterations do $\triangleright$ Step B: Policy training
13:         obtain off-policy trajectories $\hat{\tau}_{1},\ldots,\hat{\tau}_{M}$ from $\hat{x}_{1},\ldots,\hat{x}_{M}$ given by $\delta$-CS ($\mathcal{D}_{t-1},M,\delta$).
14:         obtain offline trajectory $\tau_{1},\ldots,\tau_{M}$ from $x_{1},\ldots,x_{M}\sim P_{\mathcal{D}_{t-1}}(\tau)$.
15:         train $\theta$ with TB loss over $\hat{\tau}_{1},\ldots,\hat{\tau}_{M}$ and $\tau_{1},\ldots,\tau_{M}$
$\frac{1}{2M}\sum_{i=1}^{M}\left(\log\frac{Z_{\theta}P_{F}(\tau_{i};\theta)}{R(x_{i};\phi)}\right)^{2}+\frac{1}{2M}\sum_{i=1}^{M}\left(\log\frac{Z_{\theta}P_{F}(\hat{\tau}_{i};\theta)}{R(\hat{x}_{i};\phi)}\right)^{2}.$
16:     end while
17:     obtain query samples $\hat{x}_{1},\ldots,\hat{x}_{B}$ from $\delta$-CS ($\mathcal{D}_{t-1}$, $B$, $\delta$).
18:     $\mathcal{D}_{t}\leftarrow\mathcal{D}_{t-1}\cup\{(\hat{x}_{i},f(\hat{x}_{i}))\}_{i=1}^{B}$. $\triangleright$ Step C: Dataset augmentation with oracle $f$ query
19:end for
4.1 Adjusting conservativeness parameter $\delta$
Determining the conservative parameter $\delta$ is a crucial aspect of the algorithm. We propose and study two variants, constant and adaptive $\delta$.
Constant.
As a simple approach, we set $\delta$ as a constant, selecting it to have the noise policy mask 4–15 tokens per sequence. Despite its simplicity, this choice effectively enhances policy training and leads to the discovery of high-scoring sequences during active rounds; we provide further studies on $\delta$-CS with a constant $\delta$ in Appendix B.
Adaptive.
Another intuitive approach is to adjust $\delta$ based on the uncertainty of the proxy $\sigma$ on each sequence $x$, that is $\delta(x;\sigma)$. Specifically, we define a function that assigns lower $\delta$ values for highly uncertain samples and vice versa: $\delta(x;\sigma)=\delta_{\text{const}}-\lambda\sigma(x)$. We estimate $\sigma(x)$, the standard deviation of the proxy model $f_{\phi}(x)$, via MC dropout (Gal & Ghahramani, 2016) or an ensemble method (Lakshminarayanan et al., 2017). $\lambda$ is a scaling factor and related to the influence of the proxy uncertainty on $\delta$; we set it to satisfy $\lambda\sigma\approx\frac{1}{L}$ based on the observations from the initial round.
In our main experiments, we use adaptive $\delta(x,\sigma)$ as the default setup.

### v2

@@ Abstract
Designing biological sequences with desired properties is challenging due to vast search spaces and limited evaluation budgets. Although reinforcement learning methods use proxy models for rapid reward evaluation, insufficient training data can cause proxy misspecification on out-of-distribution inputs. To address this, we propose a novel off-policy search, $\delta$-Conservative Search, that enhances robustness by restricting policy exploration to reliable regions. Starting from high-score offline sequences, we inject noise by randomly masking tokens with probability $\delta$, then denoise them using our policy. We further adapt $\delta$ based on proxy uncertainty on each data point, aligning the level of conservativeness with model confidence. Experimental results show that our conservative search consistently enhances the off-policy training, outperforming existing machine learning methods in discovering high-score sequences across diverse tasks, including DNA, RNA, protein, and peptide design.
Machine Learning, ICML
@@ 1 Introduction
Designing biological sequences with desired properties is crucial in therapeutics and biotechnology (Zimmer, 2002; Lorenz et al., 2011; Barrera et al., 2016; Sample et al., 2019; Ogden et al., 2019). However, this task is challenging due to the combinatorially large search space and the expensive and black-box nature of objective functions. Recent advances in deep learning methods for biological sequence design have shown significant promise at overcoming these challenges (Brookes & Listgarten, 2018; Brookes et al., 2019; Angermueller et al., 2020; Jain et al., 2022).
Among various approaches, reinforcement learning (RL) has emerged as one of the successful paradigms for automatic biological sequence design (Angermueller et al., 2020; Jain et al., 2022). Particularly, they employ deep neural networks as inexpensive proxy models for rapid reward evaluation by approximating expensive oracle score functions. There are two approaches for training the policy in this context: on-policy and off-policy.
@@
DyNA PPO (Angermueller et al., 2020), a representative on-policy RL method for biological sequence design, employs Proximal Policy Optimization (PPO; Schulman et al., 2017) within a proxy-based active training loop. While DyNA PPO has demonstrated effectiveness in various biological sequence design tasks, its major limitation is the limited search flexibility inherent to on-policy methods. It cannot effectively leverage offline data points, even like data collected from previous rounds.
@@
On the other hand, Generative Flow Networks (GFlowNets; Bengio et al., 2021), off-policy RL methods akin to maximum entropy policies (Tiapkin et al., 2024; Deleu et al., 2024), offer diversity-seeking capabilities and flexible exploration strategies. Jain et al. (2022) applied GFlowNets to biological sequence design with additional Bayesian active learning schemes. They leveraged the off-policy nature of GFlowNets by mixing offline datasets with on-policy data during training. This approach provided more stable training compared to DyNA PPO, resulting in better performance.
@@
However, recent studies have consistently reported that GFlowNets perform poorly in large-scale settings, such as green fluorescent protein design (Kim et al., 2023; Surana et al., 2024). We hypothesize that this bounded performance stems from the insufficient quality of the proxy model in the early rounds. Although GFlowNets can generate novel sequences beyond given data points, the resulting out-of-distribution inputs lead to unreliable rewards from the undertrained proxy (Trabucco et al., 2021; Yu et al., 2021). This motivates us to introduce a conservative search strategy to restrict the search space to the neighborhoods of the data points observed during RL training when generating sequences to query for the next round.
@@
Contribution. In this paper, we propose a novel off-policy search method called $\delta$-Conservative Search ($\delta$-CS), which enables a trade-off between sequence novelty and robustness to proxy misspecification with a conservativeness parameter $\delta$. Specifically, we iteratively train a GFlowNet using $\delta$-CS as follows: (1) we inject noise by independently masking tokens in high-score offline sequences with a Bernoulli distribution with parameter $\delta$; (2) the GFlowNet policy sequentially denoises the masked tokens; (3) we use these denoised sequences to train the policy. Figure 1 illustrates the overall procedure of $\delta$-CS. We adjust the conservative level with adaptive $\delta(x;\sigma)$ using the proxy model’s uncertainty estimates $\sigma(x)$ for each data point $x$.
@@
Our extensive experiments demonstrate that $\delta$-CS significantly improves GFlowNets, successfully discovering higher-score sequences compared to existing model-based optimization methods on diverse tasks, including DNA, RNA, protein, and peptide design. This suggests that $\delta$-CS is a robust and scalable framework for advancing research and applications in biotechnology and synthetic biology.
@@ 2 Problem Formulation
We aim to discover sequences $x\in\mathcal{V}^{L}$ that exhibit desired properties, where $\mathcal{V}$ denotes the vocabulary, such as amino acids or nucleotides, and $L$ represents the sequence length, which is usually fixed. The desired properties, such as binding affinity or enzymatic activity, are evaluated by a black-box oracle function $f:\mathcal{V}^{L}\to\mathbb{R}$. Evaluating $f$ is often both expensive and time-consuming since it typically involves wet-lab experiments or high-fidelity simulations.
Advancements in experimental techniques have enabled the parallel synthesis and evaluation of sequences in batches. Therefore, lab-in-the-loop processes are emerging as practical settings that enable active learning. Following this paradigm, we perform $T$ rounds of batch optimization, where in each round, we have the opportunity to query $B$ batched sequences to the (assumed) oracle objective function $f$. Following Angermueller et al. (2020) and Jain et al. (2022), we assume the availability of an initial offline dataset $\mathcal{D}_{0}=\{(x^{(n)},y^{(n)})\}_{n=1}^{N_{0}}$, where $y=f(x)$. The initial number of data points $N_{0}$ is typically many orders of magnitude smaller than the size of the search space, as mentioned in the introduction. The goal is to discover, after $T$ rounds, a set of sequences that are novel, diverse, and have high oracle function values.
@@ 3 Active Learning for Biological Sequence Design
Following Jain et al. (2022), we formulate an active learning process constrained by a budget of $T$ rounds with query size $B$. The active learning is conducted through an iterative procedure consisting of three standard stages, two of which are modified by our proposed method, $\delta$-Conservative Search ($\delta$-CS), which will be detailed in Section 4.
@@ 3.1 Step A: Proxy Training
Following Jain et al. (2022), we train the proxy model $f_{\phi}$ using the dataset $\mathcal{D}_{t-1}$ by minimizing the mean squared error loss:
@@ 3.2 Step B: Policy Training with $\delta$-CS
For policy training, we employ GFlowNets, [^1] Our focus is primarily on the use of GFlowNets as an off-policy reinforcement learning framework; see Section 5.2 and Section B.2 for further discussion. which aim to produce samples from a generative policy where the probability of generating a sequence $x$ is proportional to its reward, i.e.,
$p(x;\theta)\propto R(x;\phi)=\mathcal{F}(x;\phi)$
Here, $\mathcal{F}(x;\phi)$ is an acquisition function. The proxy $f_{\phi}(x)$ can be directly used as rewards, but we mainly employ the upper confidence bound (UCB; Srinivas et al., 2010) acquisition function following Jain et al. (2022).
Policy parameterization. The forward policy $P_{F}$ generates state transitions sequentially within trajectories $\tau=(s_{0}\rightarrow\ldots\rightarrow s_{L}=x)$, where $s_{0}=()$ represents the empty sequence, and each state transition involves adding a sequence token. The full sequence $s_{L}=x$ is obtained after $L$ steps, where $L$ is the sequence length. The forward policy $P_{F}(\tau;\theta)$ is a compositional policy defined as
@@
Learning objective and training trajectories. The policy is trained to minimize TB loss as follows.
@@
Usually, GFlowNets training is employed to minimize TB loss with training trajectories $\tau$ on full supports, asymptotically guaranteeing optimality for the distribution of $p(x;\theta)\propto R(x;\phi)$. A key challenge in prior works Jain et al. (2022) is that the proxy model $f_{\phi}(x)$ often produces highly unreliable rewards $R(x;\phi)$ for out-of-distribution inputs. In our approach, we mitigate this by providing off-policy trajectories within more reliable regions by injecting conservativeness into off-policy search. Therefore, we minimize TB loss with $\delta$-CS, which offers controllable conservativeness.
4 Controllable Conservativeness in Off-Policy Search
4.1 $\delta$-Conservative Search
This section introduces $\delta$-Conservative Search ($\delta$-CS), an off-policy search method that enables controllable exploration through a conservative parameter $\delta$. Here, $\delta$ defines the Bernoulli distribution governing the masking of tokens in a sequence. Our algorithm proceeds as follows:
@@ 4.1 $\delta$-Conservative Search
Rank-based reweighted prior. First, we sample a reference sequence $x$ from the prior distribution $P_{\mathcal{D}_{t-1}}$. To exploit high-scoring sequences, we employ rank-based prioritization (Tripp et al., 2020):
@@
$P_{\text{noise}}(\tilde{x}\mid x,\delta)\\ \qquad=\prod_{i=1}^{L}\left[\delta\cdot\mathbb{I}\{\tilde{e}_{i}=\text{[MASK]}\}+(1-\delta)\cdot\mathbb{I}\{\tilde{e}_{i}=e_{i}\}\right],$
@@
$P_{\text{denoise}}(\hat{e}_{t}\mid\hat{s}_{t-1},\tilde{x};\theta)\\ \qquad=\begin{cases}\mathbb{I}\{\hat{e}_{t}=\tilde{e}_{t}\},&\text{if }\tilde{e}_{t}\neq\text{[MASK]},\\ P_{F}(\hat{s}_{t}=(\hat{s}_{t-1},\hat{e}_{t})\mid\hat{s}_{t-1};\theta),&\text{if }\tilde{e}_{t}=\text{[MASK]}.\end{cases}$
@@
4.2 Adjusting Conservativeness
Determining the conservativeness parameter $\delta$ is crucial. This work proposes a simple but effective way to adjust $\delta$ based on the uncertainty of the proxy on each sequence $x$. The key insight is that even though the proxy predicts the score inaccurately, its uncertainty, which measures how the proxy gives predictions inconsistently, can guide how conservatively we search. Specifically, we define a function that assigns lower $\delta$ values for highly uncertain samples and vice versa: $\delta(x;\sigma)=\delta_{\text{const.}}-\lambda\sigma(x)$, where $\lambda$ is a scaling factor. We clamp the result to ensure it fall in $[0,1]$.
In this adaptive $\delta$ control, the reduced $\delta$ leads to a more conservative search if the proxy gives inconsistent predictions. We estimate $\sigma(x)$, the standard deviation of the proxy model $f_{\phi}(x)$, via MC dropout (Gal & Ghahramani, 2016) or an ensemble method (Lakshminarayanan et al., 2017). Note that we measure the uncertainty on the observed data points. In practice, $\delta_{\text{const.}}$ is set as 0.5 for short sequences and 0.05 for longer ones, such as proteins, as $\delta L$ tokens are masked on average. While more tailored choices of $\delta_{\text{const.}}$ could be made with task-specific knowledge, our empirical findings indicate that these simple defaults work well across a variety of settings; see Section 6.